1286 ASSUME YOU HAVE $165,000 INVESTED IN A STOCK THATS RETURNING 11.50%, $85,000 INVESTED IN A STOCK THAT IS RETURNING 22.75% & $235,000 INVESTED IN A STOCK THATS RETURNING 10.25%. WHATS THE EXPECTED RETURN OF YOUR PORTFOLIO?
A. 10.0%
B. 12.9%
C. 14.8%
D. 15.6%
I COME UP WITH 10.86%, BUT THATS NOT CORRECT, CAN YOU PLEASE SHOW THE WORK SO I WILL HAVE IT FOR FUTURE USE. ALSO, TELL ME HOW TO FIND THIS INFO USING MY TX INSTRUMENTS CALCULATOR, I DONT GET THIS, PLEASE HELP.

Answer by checkley75(3666)

165,000*.115=18975
85,000*.2275=19,337.5
235,000*.1025=24,087.5 NOW THE TOTAL INVESTMENT IS 165,000+85,000+235,000=485,000 AND THE RETURS ARE
18,975+19,337.5+24,087.5=62,400 THUS THE OVERALL PERCENTAGE IS
62,400/485,000=.1286 OR
B)12.9% ANSWER